OpenStudy (anonymous):

The local theater has 3 types of seats for Broadway plays: main floor, balcony and mezzanine. Main floor tickets are $37, balcony tickets are $31 and mezzanine are $27. One night sales are $56,591. There were 99 more main floor tickets sold than balcony and mezzanine combined. The # of balcony tickets sold is 400 more than 3 times the # of mezzanine tickets sold. How many of each type of tickets were sold?

OpenStudy (anonymous):

Let x = Main floor tickets y = Balcony tickets z = Mezzanine tickets Now \[x=y+99\] \[y-400+3z\] \[37x+31y+27z=56591\] Solve for x y and z :)

OpenStudy (anonymous):

can you help with that?

OpenStudy (anonymous):

i just need the answer :/

OpenStudy (anonymous):

Okey, lemme work it out

OpenStudy (anonymous):

thanks buddyy!

OpenStudy (anonymous):

Sorry my first equiation is wrong, its x=y+z+99

OpenStudy (anonymous):

x=883 y=688 z=96

OpenStudy (anonymous):

help with another?

OpenStudy (anonymous):

Find if no solution, only one solution and if so, find the variables, or infinitely many solutions x-2y+2z=-7 6x-4y-z=-2 7x-6y+z=6